### v1

# Inverse Purcell Suppression of Decoherence in Majorana Qubits via Environmental Engineering

Abstract
We propose a novel approach for optimizing topological quantum devices: instead of merely isolating qubits from environmental noise, we engineer the environment to actively suppress decoherence. For a Majorana qubit in a topological superconducting wire, we show that coupling via a parity-flipping operator ($\gamma_{L}$) to a broadband bosonic environment yields a decoherence rate scaling as $\Gamma\propto\rho(\epsilon)/\epsilon$, where $\rho(\epsilon)$ is the environmental density of states at the qubit splitting energy $\epsilon\sim e^{-L/\xi}$. By designing environments with suppressed density of states at low frequencies (e.g., using photonic bandgap materials or Josephson junction arrays), we achieve an “inverse Purcell effect” that leads to a suppression factor $F_{P}\propto(\epsilon/\omega_{c})^{\alpha}$, dramatically reducing decoherence. This provides a quantitative design principle where environmental engineering transforms detrimental noise into a tool for coherence stabilization. Our work establishes environmental engineering as a powerful approach for enhancing topological quantum devices.
I Introduction
The pursuit of fault-tolerant quantum computation has motivated intense research into topological qubits, with Majorana zero modes in superconducting wires emerging as a leading platform Kitaev (2001); Alicea (2012); Sau et al. (2010). Experimental progress in semiconductor-superconductor heterostructures has demonstrated signatures consistent with Majorana zero modes Mourik2012 ; Albrecht2016 , bringing topological quantum computation closer to reality. While topological protection provides inherent resilience against local perturbations, real-world implementations remain susceptible to decoherence from various environmental couplings Schlosshauer (2007), including quasiparticle poisoning Rainis2013 ; Karzig2017 and electromagnetic noise.
Traditional quantum device optimization focuses on maximizing isolation from environmental noise through improved materials and shielding. In this work, we propose and analyze a fundamentally different approach: we demonstrate that the environment itself can be engineered to actively suppress decoherence. This novel approach is motivated by recent advances in quantum circuit engineering that enable unprecedented control over electromagnetic environments Forn-Diaz2019 .
The conventional Purcell effect Purcell (1946) describes how a cavity can enhance emission rates by increasing the local density of states. Here, we demonstrate the opposite phenomenon: by engineering environments with strongly suppressed density of states at the characteristic frequency of topological qubits, we can achieve an “inverse Purcell effect” that actively protects against decoherence. This approach is particularly powerful for Majorana-based qubits, where the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ makes them naturally compatible with frequency-selective environmental engineering using existing platforms like high-impedance resonators Masluk2012 and Josephson junction arrays Pop2014 ; Manucharyan2012 ; Peltonen2013 .
In this Letter, we establish a new design principle for quantum devices: instead of treating environmental coupling as purely detrimental, we show how to harness engineered environments as a resource for coherence protection. By embedding a Majorana qubit in a carefully designed electromagnetic or phononic environment Masluk2012 ; Peltonen2013 ; Safavi-Naeini2014 with suppressed density of states at low frequencies, we transform potential decoherence pathways into tools for coherence stabilization. This approach complements recent theoretical proposals for quasiparticle-poisoning protection Karzig2017 and provides a versatile experimental pathway for enhancing topological quantum devices through environmental design.
While suppression of spontaneous emission through environmental engineering has been demonstrated in various contexts Yablonovitch1987 ; Hulet:1985zz , the application to topological quantum devices and the specific quantitative framework developed here are novel. The unique energy scales of Majorana qubits $\epsilon\sim e^{-L/\xi}$ and their parity-based encoding are specifically tailored to exploit topological protection, which makes them more effective than generic approaches.
II Model of the Majorana Qubit and Environment
We consider a topological superconducting wire of length $L$ hosting Majorana zero modes $\gamma_{L}$ and $\gamma_{R}$ at its ends. The qubit is encoded in the fermion parity basis, where the two degenerate ground states $\{|0\rangle,|1\rangle\}$ are distinguished by their total fermion parity. These states correspond to the occupation of the non-local fermion mode $f=(\gamma_{L}+i\gamma_{R})/2$, such that $|0\rangle$ is the vacuum state ($f^{\dagger}f|0\rangle=0$) and $|1\rangle$ is the occupied state ($f^{\dagger}f|1\rangle=|1\rangle$). The energy splitting $\epsilon$ between these states is exponentially small, $\epsilon\sim\Delta e^{-L/\xi}$, where $\Delta$ is the superconducting gap and $\xi$ is the coherence length; this renders them degenerate for practical purposes. This exponential suppression provides the topological protection against local perturbations that cannot simultaneously affect both widely separated Majorana modes. For our purposes, the key feature is the exponential dependence on $L/\xi$; the specific prefactor $\Delta$ can be absorbed into the overall energy scale and does not affect the functional scaling relationships that are the focus of this work. The system Hamiltonian for the isolated wire is Kitaev (2001):
$H_{\text{wire}}=\frac{i}{2}\epsilon\gamma_{L}\gamma_{R}\,.$ (1)
The decoherence mechanism arises from environmental coupling via a parity-flipping operator such as $\gamma_{L}$. While the parity operator $i\gamma_{L}\gamma_{R}$ distinguishes the parity states ($\langle 0|i\gamma_{L}\gamma_{R}|0\rangle=-1$ and $\langle 1|i\gamma_{L}\gamma_{R}|1\rangle=+1$), it is diagonal in the parity basis and cannot drive transitions between $|0\rangle$ and $|1\rangle$. However, individual Majorana operators like $\gamma_{L}$ have off-diagonal matrix elements $\langle 0|\gamma_{L}|1\rangle=1$ and $\langle 1|\gamma_{L}|0\rangle=1$, enabling them to drive transitions between the qubit states when coupled to a bosonic environment. The interaction Hamiltonian is:
$H_{\text{int}}=g\int_{0}^{L}dx\,\Phi(x)\gamma_{L}\delta(x-x_{0}),$ (2)
where $g$ is the weak coupling strength (with dimensions [M] in natural units), $\Phi$ is the bosonic field, and we approximate the localized nature of the Majorana modes by taking the coupling at a point $x_{0}$ (e.g., the center of the wire for symmetric coupling). This localized coupling approximates the dominant environmental interaction at a junction or defect site.
The parity-flipping operator $\gamma_{L}$, while built from fermionic Majorana operators, acts as a collective degree of freedom that can couple to bosonic environmental modes. This mechanism is analogous to how individual fermions in atoms or quantum dots emit photons through electronic transitions. The bosonic field $\Phi$ represents environmental modes. In the Schrödinger picture (at time $t=0$), its expansion is:
$\Phi(x)=\sum_{k}\frac{1}{\sqrt{2\omega_{k}L}}\left(b_{k}e^{ikx}+b_{k}^{\dagger}e^{-ikx}\right)\,.$ (3)
where $b_{k}^{\dagger}$ and $b_{k}$ are creation and annihilation operators for environmental bosons with frequency $\omega_{k}$. The full Hamiltonian of the combined system is then $H=H_{\text{wire}}+H_{\text{env}}+H_{\text{int}}$, where $H_{\text{env}}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the Hamiltonian of the free environment.
This localized coupling model employs a delta-function approximation $\delta(x-x_{0})$ to represent coupling concentrated at a specific location, such as a junction or defect site. This common theoretical simplification captures the essential physics while enabling analytical treatment. Although more distributed coupling would yield quantitative differences, the fundamental scaling relationship $\Gamma\propto\rho(\epsilon)/\epsilon$ derived here captures the key dependence on environmental density of states.
III Decoherence Rate Calculation
The matrix element for the transition between qubit states with emission of an environmental boson is derived using the full quantized field. For a transition from initial state $|i\rangle=|1\rangle\otimes|0_{\Phi}\rangle$ (excited qubit, vacuum environment) to final state $|f\rangle=|0\rangle\otimes|1_{k}\rangle$ (ground-state qubit, one boson in mode $k$, where $|1_{k}\rangle=b_{k}^{\dagger}|0_{\Phi}\rangle$), the matrix element is obtained from the interaction Hamiltonian (2):
$\langle f|H_{\text{int}}|i\rangle$ $=g\langle 0_{\Phi}|\langle 0|\int_{0}^{L}dx\,\Phi(x)\gamma_{L}\delta(x-x_{0})|1\rangle|0_{\Phi}\rangle$
$=g\langle 0|\gamma_{L}|1\rangle\langle 0_{\Phi}|b_{k}\Phi(x_{0})|0_{\Phi}\rangle.$ (4)
The fermionic matrix element evaluates to:
$\langle 0|\gamma_{L}|1\rangle=1\,$, (5)
since $\gamma_{L}$ acts as a parity-flipping operator that connects the two qubit states. The bosonic matrix element yields:
$\langle 0_{\Phi}|b_{k}\Phi(x_{0})|0_{\Phi}\rangle=\frac{1}{\sqrt{2\omega_{k}L}}e^{-ikx_{0}}\,.$ (6)
Thus, the complete matrix element is:
$\langle f|H_{\text{int}}|i\rangle=\frac{g}{\sqrt{2\omega_{k}L}}e^{-ikx_{0}}\,.$ (7)
The squared magnitude is:
$|\langle f|H_{\text{int}}|i\rangle|^{2}=\frac{g^{2}}{2\omega_{k}L}\,.$ (8)
The matrix element $\langle f|H_{\text{int}}|i\rangle$ is evaluated in the Schrödinger picture and is time-independent. The transition rate $\Gamma$ is given by Fermi’s Golden Rule, which involves the squared magnitude of this matrix element and ensures energy conservation.
Applying Fermi’s Golden Rule, the decoherence rate becomes:
$\Gamma=2\pi\sum_{k}|\langle f|H_{\text{int}}|i\rangle|^{2}\delta(\epsilon-\omega_{k})=\pi\sum_{k}\frac{g^{2}}{\omega_{k}L}\delta(\epsilon-\omega_{k})\,.$ (9)
Converting the sum to an integral over density of states $\rho(\omega)$, we obtain:
$\Gamma=\frac{\pi g^{2}}{L}\int\frac{\rho(\omega)}{\omega}\delta(\epsilon-\omega)\,d\omega=\frac{\pi g^{2}}{L}\frac{\rho(\epsilon)}{\epsilon}\,.$ (10)
For a 1D environment with linear dispersion $\omega_{k}=v|k|$, the free-space density of states is $\rho_{\text{free}}(\omega)=L/(\pi v)$, yielding:
$\Gamma_{\text{free}}=\frac{g^{2}}{v\epsilon}\,.$ (11)
This result reveals the fundamental challenge: for topological qubits with exponentially small $\epsilon\sim e^{-L/\xi}$, the decoherence rate diverges as $\Gamma\propto 1/\epsilon$ in a standard environment. Environmental engineering must therefore aim to suppress $\rho(\epsilon)$ rather than enhance it.
This calculation assumes weak coupling where first-order perturbation theory (Fermi’s Golden Rule) is valid, which is appropriate for the parameter regimes considered here.
IV Broadband Purcell Suppression for Coherence Enhancement
The true potential of environmental engineering emerges when we consider environments with engineered density of states. For a photonic bandgap material or specially designed cavity, we can create a density of states that is strongly suppressed at low frequencies:
[…]
where $\omega_{c}$ is a cutoff frequency, $\alpha>0$ determines the suppression strength, and $\rho_{\text{free}}(\omega)=L/(\pi v)$ is the free-space density of states for a 1D environment with linear dispersion. In the following calculation, we assume $\epsilon<\omega_{c}$ for the qubit energy splitting, which is typical for topological qubits.
The engineered decoherence rate, found by collecting Eq. (10), is:
$\Gamma_{\text{engineered}}=\frac{\pi g^{2}}{L}\frac{\rho_{\text{engineered}}(\epsilon)}{\epsilon}\,.$ (13)
Substituting $\rho_{\text{engineered}}(\epsilon)=\rho_{\text{free}}(\epsilon)(\epsilon/\omega_{c})^{\alpha}$ and $\rho_{\text{free}}(\epsilon)=L/(\pi v)$, we obtain:
$\Gamma_{\text{engineered}}=\frac{g^{2}}{v}\epsilon^{\alpha-1}\omega_{c}^{-\alpha}\,.$ (14)
Since $\epsilon\sim e^{-L/\xi}$, the $L$-dependence scales as:
$\Gamma_{\text{engineered}}\propto\epsilon^{\alpha-1}\propto e^{-(\alpha-1)L/\xi}\,.$ (15)
This scaling reveals a crucial insight: if $\alpha>1$ (strong suppression of density of states at low frequencies), then $\Gamma_{\text{engineered}}$ decreases exponentially with $L$, meaning that the coherence time increases with wire length. This means longer wires can have better coherence protection—a counterintuitive result that aligns with topological protection. If $\alpha=1$, $\Gamma_{\text{engineered}}$ is independent of $L$. If $\alpha<1$, $\Gamma_{\text{engineered}}$ increases with $L$, but this case is less desirable.
This underscores the power of environmental engineering: by choosing materials or structures with large $\alpha$ (e.g., photonic bandgaps with sharp cutoffs), we can make decoherence rates decrease with wire length, enhancing topological protection.
The resulting suppression factor compared to free space, found by collecting Eqs. (11) and (14), is:
$F_{P}=\frac{\Gamma_{\text{engineered}}}{\Gamma_{\text{free}}}=\frac{\rho_{\text{engineered}}(\epsilon)}{\rho_{\text{free}}(\epsilon)}=(\epsilon/\omega_{c})^{\alpha}\,.$ (16)
[…]
$F_{P}\sim(0.1)^{\alpha}\ll 1\,.$ (17)
providing dramatic suppression of decoherence. This “inverse Purcell” effect transforms environmental coupling from a decoherence liability into a coherence-stabilizing asset.
V Experimental Implementation and Feasibility
The environmental engineering approach proposed here is experimentally feasible with current quantum device technologies. The key requirement is creating environments with strongly suppressed density of states at low frequencies, characterized by the parameters $\omega_{c}$ (cutoff frequency) and $\alpha$ (suppression strength) introduced in Eq. (12). Several existing platforms—including high-impedance resonators, Josephson junction arrays, and phononic crystals—provide natural pathways for implementing this density of states engineering.
Superconducting microwave resonators with high characteristic impedance Forn-Diaz2019 ; Masluk2012 offer a promising platform for electromagnetic environmental engineering. These resonators exhibit modified photon mode structures that naturally suppress the density of states below a characteristic frequency $\omega_{c}\sim 1-10$ GHz, with suppression exponents $\alpha$ that can be engineered through circuit design. When coupled to semiconductor-superconductor heterostructures hosting Majorana zero modes, they can create the conditions for inverse Purcell suppression where $\rho(\epsilon)\ll\rho_{\text{free}}(\epsilon)$ for typical qubit splittings $\epsilon\sim 0.1-1$ GHz.
Josephson junction arrays provide exceptional tunability for creating bandgaps with precise control over both $\omega_{c}$ and $\alpha$ Manucharyan2012 ; Peltonen2013 . These arrays can be engineered to produce sharp cutoffs ($\alpha\gtrsim 2$) at frequencies tunable across the 0.1-10 GHz range, allowing optimal matching to specific qubit parameters. The ability to dynamically tune $\omega_{c}$ in situ through external flux biases makes this platform particularly attractive for verifying the predicted scaling for suppression factor $F_{P}\propto(\epsilon/\omega_{c})^{\alpha}$.
For phonon-mediated decoherence channels, phononic crystals Safavi-Naeini2014 can create bandgaps that suppress acoustic modes at relevant frequencies. By patterning substrates with periodic structures having lattice constants $a\sim v_{s}/\omega_{c}$ (where $v_{s}$ is the sound velocity), bandgaps can be created at $\omega_{c}\sim 1-10$ GHz. Phononic crystal engineering has already produced quality factors $Q>10^{6}$ in mechanical resonators Safavi-Naeini2014 , demonstrating the feasibility of strong suppression of the phononic density of states $\rho(\omega)$ around a target frequency. Within our theoretical framework, such suppression applied to the electromagnetic or phononic environment of a superconducting qubit would create the conditions needed for the inverse Purcell effect, dramatically reducing decoherence rates.
The required parameter regime $\omega_{c}\sim 1-10$ GHz and $\alpha>1$ is readily achievable in these platforms. For a topological qubit with $\epsilon\sim 1$ GHz and an environment with $\omega_{c}\sim 10$ GHz and $\alpha=2$, the predicted suppression factor is $F_{P}\sim(0.1)^{2}=0.01$, representing a 100-fold reduction in decoherence rate.
These implementation pathways are not mutually exclusive; hybrid approaches combining electromagnetic and phononic engineering may provide comprehensive protection against multiple decoherence channels. The flexibility of environmental engineering allows customization for specific device architectures, making it a versatile tool for optimizing topological quantum devices.
VI Conclusion and Experimental Implications
Our analysis demonstrates that environmental engineering provides a powerful and unconventional mechanism for controlling coherence in topological quantum devices. Rather than attempting to isolate qubits from environmental noise, we have shown that strategic suppression of the environmental density of states at low frequencies can dramatically reduce decoherence rates through an inverse Purcell effect. The derived scaling relation $\Gamma=\frac{\pi g^{2}}{L}\frac{\rho(\epsilon)}{\epsilon}$ with $\epsilon\sim e^{-L/\xi}$ establishes a quantitative design principle: coherence times can be optimized by engineering environments with strongly suppressed density of states at the characteristic energy scales of topological qubits.
This approach is particularly well-suited for Majorana-based qubits, where the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ naturally places the relevant energy scales in the microwave regime (0.1-10 GHz). This aligns perfectly with the operational range of superconducting quantum circuits, and recent advances in high-impedance resonators, Josephson junction arrays, and phononic crystals provide precisely the tools needed for implementing the required density of states engineering.
Our work establishes environmental engineering as a powerful complementary approach to intrinsic topological protection. The key insight is that environmental coupling can be transformed from a limitation into a resource when properly engineered. By designing environments with suppressed density of states at low frequencies, we can achieve dramatic suppression of decoherence rates, with suppression factors reaching $F_{P}\sim(0.1)^{\alpha}\ll 1$ for realistic parameters.
These findings open new avenues for optimizing topological quantum devices through deliberate electromagnetic and phononic design. By engineering environments with specific density of states profiles, we can potentially achieve unprecedented control over quantum coherence in non-inertial frames, advancing the pursuit of fault-tolerant quantum computation.
Beyond topological quantum computing, this environmental engineering approach could find applications in other quantum systems, such as quantum optics and cavity QED, where control over the electromagnetic environment is crucial, thus establishing environmental design as a complementary tool alongside error correction.