### v2

# Inverse Purcell Suppression of Decoherence in Majorana Qubits via Environmental Engineering

Abstract
We propose a novel approach for optimizing topological quantum devices: instead of merely isolating qubits from environmental noise, we engineer the environment to actively suppress decoherence. For a Majorana qubit in a topological superconducting wire, the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ provides protection against local perturbations but renders it highly susceptible to pure dephasing from low-frequency environmental noise. We show that coupling via a parity-conserving operator ($i\gamma_{L}\gamma_{R}$) to a bosonic environment yields a dephasing rate $\Gamma_{\phi}\propto S(\epsilon)$, where $S(\epsilon)$ is the environmental noise power at the qubit splitting frequency. In the experimentally relevant regime where $k_{B}T\gtrsim\hbar\epsilon$ (with $T\sim 10-100$ mK), the noise power scales as $S(\epsilon)\propto\rho(\epsilon)k_{B}T/\hbar\epsilon$, leading to a dephasing rate $\Gamma_{\phi}\propto\rho(\epsilon)T/\epsilon$. This exposes a fundamental challenge: the dephasing rate diverges as $1/\epsilon$ for a standard environment, e.g., a 1D system with linear dispersion where $\rho(\epsilon)$ is constant. We overcome this by designing environments with a suppressed density of states following $\rho_{\text{engineered}}(\epsilon)=\rho_{\text{free}}(\epsilon)(\epsilon/\omega_{c})^{\alpha}$. This creates an “inverse Purcell effect” that yields a temperature-independent suppression factor $F_{P}=(\epsilon/\omega_{c})^{\alpha}$. For $\alpha>1$, the engineered dephasing rate decreases exponentially with wire length, $\Gamma_{\phi,\text{engineered}}\propto e^{-(\alpha-1)L/\xi}$, meaning longer wires provide better coherence protection. This provides a quantitative design principle where environmental engineering transforms detrimental noise into a tool for coherence stabilization, while respecting fermion parity superselection rules. Our work establishes environmental engineering as a powerful approach for enhancing topological quantum devices.
I Introduction
The pursuit of fault-tolerant quantum computation has motivated intense research into topological qubits, with Majorana zero modes in superconducting wires emerging as a leading platform Kitaev2001 ; Alicea2012 ; Sau2010 . Experimental progress in semiconductor-superconductor heterostructures has demonstrated signatures consistent with Majorana zero modes Mourik2012 ; Albrecht2016 , bringing topological quantum computation closer to reality. While topological protection provides inherent resilience against local perturbations through exponentially small energy splittings $\epsilon\sim e^{-L/\xi}$, this very property creates a fundamental vulnerability: such small splittings render the qubits highly susceptible to pure dephasing from low-frequency environmental noise. Real-world implementations therefore remain susceptible to decoherence from various environmental couplings Schlosshauer2007 , including quasiparticle poisoning Rainis2013 ; Karzig2017 (a relaxation error) and broadband electromagnetic noise (a primary source of dephasing).
Traditional quantum device optimization focuses on maximizing isolation from environmental noise through improved materials and shielding. In this work, we propose and analyze a fundamentally different approach: we demonstrate that the environment itself can be engineered to actively suppress decoherence by targeting this specific vulnerability. This novel approach is motivated by recent advances in quantum circuit engineering that enable unprecedented control over electromagnetic environments Forn-Diaz2019 .
The conventional Purcell effect Purcell1946 describes how a cavity can enhance emission and relaxation rates by increasing the local density of states. Here, we demonstrate an analogous but opposite phenomenon for dephasing: by engineering environments with strongly suppressed density of states at the specific frequencies relevant to the qubit’s energy splitting, we achieve an “inverse Purcell effect” that actively protects against phase randomization. This approach is particularly powerful for Majorana-based qubits, where the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ makes them naturally compatible with frequency-selective environmental engineering. Furthermore, their parity-based encoding requires coupling via a parity-conserving operator ($i\gamma_{L}\gamma_{R}$), which intrinsically suppresses relaxation and isolates dephasing as the dominant decoherence channel to target. Existing platforms like high-impedance resonators Masluk2012 and Josephson junction arrays Pop2014 ; Manucharyan2012 ; Peltonen2013 provide the necessary tools to implement this spectral shaping.
In this Letter, we establish a new design principle for quantum devices: instead of treating environmental coupling as purely detrimental, we show how to harness engineered environments as a resource for coherence protection. By embedding a Majorana qubit in a carefully designed electromagnetic or phononic environment Masluk2012 ; Peltonen2013 ; Safavi-Naeini2014 with a suppressed density of states around its transition frequency $\epsilon$, we transform the vulnerability to low-frequency noise into a tool for coherence stabilization. Specifically, we demonstrate that proper environmental engineering can yield an exponential suppression of dephasing rates with increasing wire length ($\Gamma_{\phi}\propto e^{-(\alpha-1)L/\xi}$) and create temperature-independent protection. This approach complements recent theoretical proposals for quasiparticle-poisoning protection Karzig2017 and provides a versatile experimental pathway for enhancing topological quantum devices through environmental design.
While suppression of spontaneous emission through environmental engineering has been demonstrated in various contexts Yablonovitch1987 ; Hulet:1985zz , the application to topological quantum devices and the specific quantitative framework we develop for dephasing suppression are novel. The unique energy scales of Majorana qubits $\epsilon\sim e^{-L/\xi}$ and their parity-based encoding are specifically tailored to exploit this form of protection, making the environmental engineering approach more effective for them than for generic qubit systems.
II Model of the Majorana Qubit and Environment
We consider a topological superconducting wire of length $L$ hosting Majorana zero modes $\gamma_{L}$ and $\gamma_{R}$ at its ends. The qubit is encoded in the fermion parity basis, where the two degenerate ground states $\{|0\rangle,|1\rangle\}$ are distinguished by their total fermion parity. These states correspond to the occupation of the non-local fermion mode $f=(\gamma_{L}+i\gamma_{R})/2$, such that $|0\rangle$ is the vacuum state ($f^{\dagger}f|0\rangle=0$) and $|1\rangle$ is the occupied state ($f^{\dagger}f|1\rangle=|1\rangle$). The energy splitting $\epsilon$ between these states is exponentially small, $\epsilon\sim\Delta e^{-L/\xi}$, where $\Delta$ is the superconducting gap and $\xi$ is the coherence length; this renders them degenerate for practical purposes. This exponential suppression provides the topological protection against local perturbations that cannot simultaneously affect both widely separated Majorana modes. However, this very smallness creates a fundamental vulnerability: the qubit operates at microwave frequencies ($\epsilon\sim 0.1-10$ GHz) where environmental noise is significant, making it highly susceptible to dephasing. For our purposes, the key feature is the exponential dependence on $L/\xi$; the specific prefactor $\Delta$ can be absorbed into the overall energy scale and does not affect the functional scaling relationships that are the focus of this work. The system Hamiltonian for the isolated wire is Kitaev2001 :
$H_{\text{wire}}=\frac{i}{2}\epsilon\gamma_{L}\gamma_{R}\,.$ (1)
The decoherence mechanism arises specifically from environmental coupling that causes pure dephasing. We use a parity-conserving operator $i\gamma_{L}\gamma_{R}$ that commutes with the total fermion parity and thus cannot cause transitions between parity sectors; instead, it modulates the energy splitting between states $|0\rangle$ and $|1\rangle$, leading exclusively to dephasing. The interaction Hamiltonian is:
$H_{\text{int}}=g\int_{0}^{L}dx\,\Phi(x)(i\gamma_{L}\gamma_{R})\delta(x-x_{0})\,$, (2)
where $g$ is the weak coupling strength, $\Phi$ is the bosonic field, and we approximate localized coupling at a point $x_{0}$ (e.g., the center of the wire for symmetric coupling). This choice respects the fermion parity superselection rule Wick:1970bd ; Hassler2014 , as $i\gamma_{L}\gamma_{R}$ is a bosonic operator built from an even number of fermionic operators.
The bosonic field $\Phi$ represents environmental modes that will cause dephasing through their low-frequency components. In the Schrödinger picture (at time $t=0$), its expansion is:
$\Phi(x)=\sum_{k}\frac{1}{\sqrt{2\omega_{k}L}}\left(b_{k}e^{ikx}+b_{k}^{\dagger}e^{-ikx}\right),$ (3)
where $b_{k}^{\dagger}$ and $b_{k}$ are creation and annihilation operators for environmental bosons with frequency $\omega_{k}$. For a 1D environment with linear dispersion $\omega_{k}=v|k|$, the density of states per unit frequency is constant, $\rho(\omega)=L/(\pi v)$. This density of states directly determines the noise power spectrum $S(\omega)$ that governs the dephasing rate, as will be shown in the next section. The full Hamiltonian of the combined system is then $H=H_{\text{wire}}+H_{\text{env}}+H_{\text{int}}$, where $H_{\text{env}}=\sum_{k}\omega_{k}b_{k}^{\dagger}b_{k}$ is the Hamiltonian of the free environment.
This localized coupling model employs a delta-function approximation $\delta(x-x_{0})$ to represent coupling concentrated at a specific location, such as a junction or defect site. This common theoretical simplification captures the essential physics while enabling analytical treatment. Although more distributed coupling would yield quantitative differences, the fundamental scaling relationship $\Gamma_{\phi}\propto S(\epsilon)$ derived here captures the key dependence on environmental noise power, particularly the critical role of low-frequency noise due to the small energy splitting $\epsilon$.
III Dephasing Rate Calculation
The dephasing rate for a qubit with energy splitting $\epsilon$ is determined by the noise power spectrum $S(\omega)$ of the environment. For the parity-conserving interaction $H_{\text{int}}=g(i\gamma_{L}\gamma_{R})\Phi$, which commutes with the system Hamiltonian, the dephasing rate follows the standard result for pure dephasing Schoelkopf2003 ; Clerk2010 :
$\Gamma_{\phi}=\frac{g^{2}}{2}S(\epsilon),$ (4)
where $S(\epsilon)$ is the noise power at frequency $\epsilon$. This expression derives from the mode expansion of $\Phi$ and the fluctuation-dissipation theorem Callen1951 ; Kubo1966 , which relates $S(\omega)$ to the density of states $\rho(\omega)$ and temperature. For a bath of harmonic oscillators, this yields Clerk2010 ; Gardiner2004 :
$S(\omega)=2\pi\rho(\omega)\left[n(\omega)+\frac{1}{2}\right]=\pi\rho(\omega)\coth\left(\frac{\hbar\omega}{2k_{B}T}\right),$ (5)
where $n(\omega)=(e^{\hbar\omega/k_{B}T}-1)^{-1}$ is the Bose-Einstein distribution. In the low temperature limit ( $k_{B}T\ll\hbar\omega$), the noise power simplifies to $S(\omega)\approx\pi\rho(\omega)$, while in the high-temperature limit ($k_{B}T\gg\hbar\omega$), it scales as $S(\omega)\approx 2\pi\rho(\omega)\frac{k_{B}T}{\hbar\omega}$.
For Majorana qubits with $\epsilon\sim 1$ GHz ($\hbar\epsilon/k_{B}\sim 48$ mK) and operating at temperatures $T\sim 10-100$ mK, we typically satisfy $k_{B}T\gtrsim\hbar\epsilon$. This places us in the regime where the high-temperature approximation applies:
$S(\epsilon)\approx 2\pi\rho(\epsilon)\frac{k_{B}T}{\hbar\epsilon}\,$, (6)
and consequently, the dephasing rate becomes:
$\Gamma_{\phi,\text{free}}\approx\frac{\pi g^{2}k_{B}T}{\hbar\epsilon}\rho_{\text{free}}(\epsilon)\,.$ (7)
This result reveals the fundamental challenge: for topological qubits with exponentially small $\epsilon\sim e^{-L/\xi}$, the dephasing rate in a standard environment diverges as $\Gamma_{\phi,\text{free}}\propto 1/\epsilon$ because the thermal factor $k_{B}T/(\hbar\epsilon)$ grows as $1/\epsilon$ while $\rho_{\text{free}}(\epsilon)$ remains approximately constant for a 1D environment with linear dispersion $\omega_{k}=v|k|$. Environmental engineering must therefore aim to suppress $\rho(\epsilon)$ rather than enhance it to counteract this divergence.
IV Broadband Purcell Suppression for Coherence Enhancement
The true potential of environmental engineering emerges when we consider environments with engineered density of states. For a photonic bandgap material or specially designed cavity, we can create a density of states that is strongly suppressed at low frequencies:
$\frac{\rho_{\text{engineered}}(\omega)}{\rho_{\text{free}}(\omega)}=\begin{cases}(\omega/\omega_{c})^{\alpha},&\omega<\omega_{c},\\ 1,&\omega\geq\omega_{c},\end{cases}$ (8)
where $\omega_{c}$ is a cutoff frequency, $\alpha>0$ determines the suppression strength, and $\rho_{\text{free}}(\omega)=L/(\pi v)$ is the free-space density of states for a 1D environment with linear dispersion. In the following calculation, we assume $\epsilon<\omega_{c}$ for the qubit energy splitting, which is typical for topological qubits.
The engineered dephasing rate is determined by the noise power spectrum at the qubit frequency $\epsilon$:
$\Gamma_{\phi,\text{engineered}}=\frac{g^{2}}{2}S(\epsilon).$ (9)
For a bosonic environment, the noise power spectrum follows from the fluctuation-dissipation theorem:
$S(\epsilon)=\pi\rho(\epsilon)\coth\left(\frac{\hbar\epsilon}{2k_{B}T}\right).$ (10)
In the experimentally relevant regime where $k_{B}T\gtrsim\hbar\epsilon$ (typical for Majorana qubits with $\epsilon\sim 1$ GHz and $T\sim 10-100$ mK), we have:
$S(\epsilon)\approx 2\pi\rho(\epsilon)\frac{k_{B}T}{\hbar\epsilon},$ (11)
leading to the engineered dephasing rate:
$\Gamma_{\phi,\text{engineered}}\approx\frac{\pi g^{2}k_{B}T}{\hbar\epsilon}\rho_{\text{engineered}}(\epsilon).$ (12)
Using the engineered spectral density, $\rho_{\text{engineered}}(\epsilon)=\rho_{\text{free}}(\epsilon)(\epsilon/\omega_{c})^{\alpha}$, and since $\rho_{\text{free}}(\epsilon)=L/(\pi v)$ is approximately constant for small $\epsilon$, we obtain:
$\Gamma_{\phi,\text{engineered}}\propto\frac{g^{2}k_{B}T}{\hbar}\epsilon^{\alpha-1}\omega_{c}^{-\alpha}.$ (13)
Since $\epsilon\sim e^{-L/\xi}$, the $L$-dependence scales as:
$\Gamma_{\phi,\text{engineered}}\propto\epsilon^{\alpha-1}\propto e^{-(\alpha-1)L/\xi}.$ (14)
This scaling reveals a crucial insight: if $\alpha>1$ (strong suppression of density of states at low frequencies), then $\Gamma_{\phi,\text{engineered}}$ decreases exponentially with $L$, meaning that the coherence time increases with wire length. This counterintuitive result—that longer wires provide better coherence protection—aligns perfectly with topological protection.
The resulting suppression factor compared to free space is:
$F_{P}=\frac{\Gamma_{\phi,\text{engineered}}}{\Gamma_{\phi,\text{free}}}=\frac{\rho_{\text{engineered}}(\epsilon)}{\rho_{\text{free}}(\epsilon)}=(\epsilon/\omega_{c})^{\alpha}.$ (15)
Remarkably, the temperature dependence cancels in this ratio, making the suppression factor $F_{P}$ temperature-independent. This is a key advantage of our approach: environmental engineering provides protection against dephasing that is robust against thermal fluctuations in the experimentally relevant regime.
[…]
$F_{P}\sim(0.1)^{\alpha}\ll 1,$ (16)
providing dramatic suppression of dephasing. This “inverse Purcell” effect transforms environmental coupling from a decoherence liability into a coherence-stabilizing asset.
V Experimental Implementation and Feasibility
The environmental engineering approach proposed here is experimentally feasible with current quantum device technologies. The key requirement is creating environments with strongly suppressed density of states at low frequencies, characterized by the parameters $\omega_{c}$ (cutoff frequency) and $\alpha$ (suppression strength). Several existing platforms—including high-impedance resonators, Josephson junction arrays, and phononic crystals—provide natural pathways for implementing this density of states engineering. Critically, these platforms operate at temperatures $T\sim 10-100$ mK, which aligns perfectly with the experimental conditions for Majorana qubit devices and ensures that the low-temperature approximation ($k_{B}T\gtrsim\hbar\epsilon$) remains valid.
Superconducting microwave resonators with high characteristic impedance Forn-Diaz2019 ; Masluk2012 offer a promising platform for electromagnetic environmental engineering. These resonators exhibit modified photon mode structures that naturally suppress the density of states below a characteristic frequency $\omega_{c}\sim 1-10$ GHz, with suppression exponents $\alpha$ that can be engineered through circuit design. When coupled to semiconductor-superconductor heterostructures hosting Majorana zero modes at temperatures $T\sim 10-100$ mK, they can create the conditions for inverse Purcell suppression where $\rho(\epsilon)\ll\rho_{\text{free}}(\epsilon)$ for typical qubit splittings $\epsilon\sim 0.1-1$ GHz. Recent experiments have demonstrated such suppression in high-impedance resonators integrated with superconducting qubits, confirming the viability of this approach Grimm2020 .
Josephson junction arrays provide exceptional tunability for creating bandgaps with precise control over both $\omega_{c}$ and $\alpha$ Manucharyan2012 ; Peltonen2013 . These arrays can be engineered to produce sharp cutoffs ($\alpha\gtrsim 2$) at frequencies tunable across the 0.1-10 GHz range, allowing optimal matching to specific qubit parameters. The ability to dynamically tune $\omega_{c}$ in situ through external flux biases makes this platform particularly attractive for verifying the predicted scaling for suppression factor $F_{P}\propto(\epsilon/\omega_{c})^{\alpha}$. Experimental studies have shown that Josephson arrays can achieve high-quality bandgaps with minimal losses at cryogenic temperatures ($T<100$ mK), as reported in recent works Pop2014 ; Rymarz2021 .
For phonon-mediated decoherence channels, phononic crystals Safavi-Naeini2014 can create bandgaps that suppress acoustic modes at relevant frequencies. By patterning substrates with periodic structures having lattice constants $a\sim v_{s}/\omega_{c}$ (where $v_{s}$ is the sound velocity), bandgaps can be created at $\omega_{c}\sim 1-10$ GHz. Phononic crystal engineering has already produced quality factors $Q>10^{6}$ in mechanical resonators Safavi-Naeini2014 , demonstrating the feasibility of strong suppression of the phononic density of states $\rho(\omega)$ around a target frequency. This experimental progress in nanofabrication and material design directly supports the feasibility of implementing our proposed environmental engineering approach for phonon noise suppression.
The required parameter regime $\omega_{c}\sim 1-10$ GHz and $\alpha>0$ is readily achievable in these platforms. For a topological qubit with $\epsilon\sim 1$ GHz and an environment with $\omega_{c}\sim 10$ GHz and $\alpha=2$, the predicted suppression factor is $F_{P}\sim(0.1)^{2}=0.01$, representing a 100-fold reduction in dephasing rate. While these platforms show great promise, challenges such as fabrication complexity, parasitic losses, and mode hybridization must be addressed in practical implementations. However, current advances in nanofabrication and materials science are rapidly mitigating these issues, making environmental engineering increasingly viable.
These implementation pathways are not mutually exclusive; hybrid approaches combining electromagnetic and phononic engineering may provide comprehensive protection against multiple decoherence channels. The flexibility of environmental engineering allows customization for specific device architectures, making it a versatile tool for optimizing topological quantum devices.
VI Conclusion and Experimental Implications
Our analysis demonstrates that environmental engineering provides a powerful and unconventional mechanism for controlling coherence in topological quantum devices. The exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ that underpins topological protection also creates a fundamental vulnerability to low-frequency noise, leading to pure dephasing as the dominant decoherence channel. Rather than attempting to isolate qubits from environmental noise, we have shown that strategic suppression of the environmental density of states at low frequencies can dramatically reduce dephasing rates through an inverse Purcell effect. The derived scaling relation $\Gamma_{\phi}\propto S(\epsilon)$ with $S(\epsilon)\propto\rho(\epsilon)\frac{k_{B}T}{\hbar\epsilon}$ in the experimentally relevant regime ($k_{B}T\gtrsim\hbar\epsilon$) establishes a quantitative design principle: coherence times can be optimized by engineering environments with strongly suppressed density of states $\rho(\epsilon)$ at the characteristic energy scales of topological qubits.
This approach is particularly well-suited for Majorana-based qubits, where the exponentially small energy splitting $\epsilon\sim e^{-L/\xi}$ naturally places the relevant energy scales in the microwave regime (0.1-10 GHz). This aligns perfectly with the operational range of existing superconducting quantum circuits. Crucially, recent experimental advances—such as high-impedance resonators Masluk2012 , tunable Josephson junction arrays Manucharyan2012 , and phononic crystals Safavi-Naeini2014 —provide precisely the tools needed to implement the required spectral engineering at millikelvin temperatures ($T\sim 10-100$ mK). These platforms directly address the vulnerability imposed by the small $\epsilon$ by enabling targeted suppression of the noise power spectrum.
Our work establishes environmental engineering as a powerful complementary approach to intrinsic topological protection. The key insight is that environmental coupling can be transformed from a limitation into a resource when properly engineered. By designing environments with a suppressed density of states following $\rho_{\text{engineered}}(\epsilon)/\rho_{\text{free}}(\epsilon)=(\epsilon/\omega_{c})^{\alpha}$, we achieve a suppression factor $F_{P}=(\epsilon/\omega_{c})^{\alpha}$ that is notably temperature-independent. This robustness against thermal fluctuations allows for dramatic suppression of dephasing, with $F_{P}\sim(0.1)^{\alpha}\ll 1$ for realistic parameters ($\epsilon\sim 1$ GHz, $\omega_{c}\sim 10$ GHz, $\alpha=2$), representing a 100-fold reduction in dephasing rate.
These findings open new avenues for optimizing topological quantum devices through deliberate electromagnetic and phononic design. The fact that the dephasing rate scales as $\Gamma_{\phi}\propto\epsilon^{\alpha-1}\propto e^{-(\alpha-1)L/\xi}$ implies that for $\alpha>1$, longer wires (with smaller $\epsilon$) inherently provide better coherence protection—a counterintuitive result that aligns perfectly with topological protection. By engineering environments with specific spectral profiles, we can potentially achieve unprecedented control over quantum coherence, advancing the pursuit of fault-tolerant quantum computation.
Beyond topological quantum computing, this environmental engineering approach could find applications in other quantum systems where low-frequency noise is a limiting factor, such as quantum optics and cavity QED Forn-Diaz2019 , thus establishing environmental design as a versatile tool alongside dynamical error correction.